OpenStudy (anonymous):

Assumed that all variables are nonzero Simplif(x/2)^7 I have solved it ,I just want to confirmed if my answer is correct

OpenStudy (anonymous):

ok will confirm when done with quiz

OpenStudy (anonymous):

1- (x/2)7= x7/27 = x7/128

OpenStudy (anonymous):

36x^2/6x